Compute the electric field experienced by a negative test charge given the following conditions in a vacuum: b. Source charge: +6.78 × 10^–15 C Distance from source charge: 4.6 × 10^–19 m (Your answer must be in this format : 1.11x10^1 N/C, -1.11x10^-1 N/C) 


1
Expert's answer
2022-03-07T10:19:08-0500
"E=k\frac{q}{r^2}\\
={\rm 9*10^9\: N\cdot m^2/C^2*\frac{6.78 × 10^{–15}\: C }{(4.6 × 10^{–19}\: m)^2}}\\
=2.88*10^{32}\:\rm V/m"